# UVIP: Model-Free Approach to Evaluate Reinforcement Learning Algorithms

### v4

@@ 7 Conclusion and future work
In this work we propose a new approach towards model-free evaluation of the agent’s policies in RL, based on upper solutions to the Bellman optimality equation (1). To the best of our knowledge, the UVIP is the first procedure which allows to construct the non-asymptotic confidence intervals for the optimal value function $V^{\star}$ based on the value function corresponding to an arbitrary policy $\pi$. In our analysis we consider only infinite-horizon MDPs and assume that sampling from the conditional distribution ${\rm P}^{a}(\cdot|x)$ is feasible for any $x\in\mathsf{X}$ and $a\in\mathsf{A}.$ A promising future research direction is to generalize UVIP to the case of finite-horizon MDPs combining it with the idea of Real-time dynamic programming (see Efroni et al. (2019)). Moreover, the plain Monte Carlo estimates are not necessarily the most efficient way to estimate the outer expectation in Algorithm $1$. Other stochastic approximation techniques could also be applied to approximate the solution of (2).
8 Acknowledgments
The article was prepared within the framework of the HSE University Basic Research Program. This research was supported in part through computational resources of HPC facilities at HSE University Kostenetskiy et al. (2021).
@@ Appendix A Proof of the main results
Throughout this section we will use additional notations. Let $\psi_{2}(x)=e^{x^{2}}-1$, $x\in\mathbb{R}$. For r.v. $\eta$ we denote $\|\eta\|_{\psi_{2}}\doteq\inf\{t>0:\mathsf{E}\left[\exp\{\eta^{2}/t^{2}\}\right]\leq 2\}$ the Orlicz 2-norm. We say that $\eta$ is a sub-Gaussian random variable if $\|\eta\|_{\psi_{2}}<\infty$. In particular, this implies that for some constants $C,c>0$, $\mathrm{P}(|\eta|\geq t)\leq 2\exp\{-ct^{2}/\|\eta\|_{\psi_{2}}^{2}\}$ and $\mathsf{E}^{1/p}[|\eta|^{p}]\leq C\sqrt{p}\|\eta\|_{\psi_{2}}$ for all $p\geq 1$. Consider a random process $(X_{t})_{t\in T}$ on a metric space $(T,\mathsf{d})$. We say that the process has sub-Gaussian increments if there exists $K\geq 0$ such that
@@
Proof.
We apply the empirical process methods. To simplify notations we denote
@@
that is, $Z(x,a)$ is a random process on the metric space $(\mathsf{X}\times\mathsf{A},\rho)$. Below we show that the process $Z(x,a)$ has sub-Gaussian increments. In order to show it, let us introduce for $\ell\in[M]$
$Z_{\ell}\doteq[V^{\pi}(\psi(x,a,\xi_{l}))-\mathsf{E}V^{\pi}(\psi(x,a,\xi_{l}))]-[V^{\pi}(\psi(x^{\prime},a^{\prime},\xi_{l}))-\mathsf{E}V^{\pi}(\psi(x^{\prime},a^{\prime},\xi_{l}))]\,.$
@@
that is, $Z_{\ell}$ is a sub-Gaussian r.v. for any $\ell\in[M]$. Since $Z(x,a)-Z(x^{\prime},a^{\prime})=M^{-1/2}\sum_{\ell=1}^{M}Z_{\ell}$ is a sum of independent sub-Gaussian r.v, we may apply (Vershynin, 2018, Proposition 2.6.1 and Eq. (2.16))) to obtain that $Z(x,a)$ has sub-Gaussian increments with parameter $K\asymp L_{\pi}$. Fix some $(x_{0},a_{0})\in\mathsf{X}\times\mathsf{A}$. By the triangular inequality,
@@
By the Dudley integral inequality, e.g. (Vershynin, 2018, Theorem 8.1.6), for any $\delta\in(0,1)$,
$\sup_{(x,a),(x^{\prime},a^{\prime})\in\mathsf{X}\times\mathsf{A}}|Z(x,a)-Z(x^{\prime},a^{\prime})|\lesssim L_{\pi}\bigl{[}I_{\mathcal{D}}+\mathsf{D}\sqrt{\log(2/\delta)}\bigr{]}\,.$
@@
holds with probability $1-\delta$. The last two inequalities and (15) imply the statement. ∎
@@
$\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)=\frac{1}{M_{2}}\sum_{j=M_{1}+1}^{M_{1}+M_{2}}\max_{a}\left\{r^{a}(x)+\gamma\Bigl{(}\widehat{V}_{k}^{{\rm{up}}}(Y_{j}^{x,a})-V^{\pi}(Y_{j}^{x,a})+\frac{1}{M_{1}}\sum_{\ell=1}^{M_{1}}V^{\pi}(Y_{\ell}^{x,a})\Bigr{)}\right\}$
@@
Using that $\gamma L_{\psi}<1$, the maximal Lipshitz constant of $\widetilde{V}_{k}^{{\rm{up},\pi}}(x),k\in\mathbb{N}$ is uniformly bounded by
@@
$\begin{split}&\mathsf{E}^{1/p}[|R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})-V^{*}(x)|^{p}]\leq\\ &\mathsf{E}^{1/p}\left[\left|\max_{a}\left\{r^{a}(x)+\gamma\Bigl{(}\widehat{V}_{k}^{{\rm{up}}}(Y_{j}^{x,a})-V^{\pi}(Y_{j}^{x,a})+M_{1}^{-1}\sum\nolimits_{\ell=1}^{M_{1}}V^{\pi}(Y_{\ell}^{x,a})\Bigr{)}\right\}\right.\right.-\\ &\qquad\left.\left.\max_{a}\left\{r^{a}(x)+\gamma(V^{\star}(Y^{x,a})-V^{\star}(Y^{x,a})+{\rm P}^{a}V^{\star}(x)\right\}\right|^{p}\right]\end{split}$
Hence, with the Minkowski inequality and $|{\rm P}^{a}V^{\star}(x)-\mathsf{E}V^{\pi}(\psi(x,a,\cdot))|\leq\|V^{\pi}-V^{*}\|_{\mathsf{X}}$, we get
$\begin{split}&\mathsf{E}^{1/p}[|R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})-V^{*}(x)|^{p}]\leq\gamma{\mathcal{M}}_{k}+2\gamma\left\|V^{\pi}-V^{{\rm*}}\right\|_{\mathsf{X}}\\ &\qquad\qquad\qquad+\gamma\mathsf{E}^{1/p}\Bigl{[}\Bigl{\|}M_{1}^{-1}\sum\nolimits_{\ell=1}^{M_{1}}[V^{\pi}(\psi(\cdot,\cdot,\xi_{k,\ell}))-\mathsf{E}V^{\pi}(\psi(\cdot,\cdot,\xi_{k,\ell}))]\Bigr{\|}_{\mathsf{X}\times\mathcal{A}}^{p}\Bigr{]}\end{split}$ (19)
To analyze the last term we use the empirical process methods. By Proposition A.1, we get
@@
Furthermore, with (7) we construct Lipshitz interpolant $\widehat{V}_{k+1}^{{\rm{up}}}$, such that
@@
Proof of Corollary 5.1.
Consider $\rho((x,a),(x^{\prime},a^{\prime}))=\mathbbm{1}_{\{(x,a)\neq(x^{\prime},a^{\prime})\}}$ and $\mathsf{X}_{N}=\mathsf{X}$, that is, we bypass the approximation step. Then $\mathsf{D}=1$, $I_{\mathcal{D}}\lesssim\sqrt{\log(|\mathsf{X}||\mathsf{A}|)}$, $\rho(\mathsf{X}_{N},\mathsf{X})=0$, and $r^{a}(\cdot)$ is Lipshitz w.r.t. $\rho_{\mathsf{X}}$ with $L_{\max}\leq R_{\max}$. Moreover, one can take $L_{\psi}=1$ and $L_{\pi}=R_{\max}/(1-\gamma)$ in Assumption A4. Hence, A1 – A4 are valid and one may apply Theorem 5.1. Bound (21) in this case writes as
@@
We use definition of $\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)$ and $R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})$ from Theorem 5.1. To simplify notations we denote $\boldsymbol{\xi}_{k,M_{1}}=(\xi_{k,1},\ldots,\xi_{k,M_{1}})$ and $\boldsymbol{\xi}_{k,M_{2}}=(\xi_{k,M_{1}+1},\ldots,\xi_{k,M_{1}+M_{2}})$. In these notations $\boldsymbol{\xi}_{k}=(\boldsymbol{\xi}_{k,M_{1}},\boldsymbol{\xi}_{k,M_{2}})$. Recall that, by the construction, $\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)$ can be evaluated only at the points $x\in\{\mathsf{x}_{1},\dots,\mathsf{x}_{N}\}$. By the definition,
@@
where $\xi$ is an i.i.d. copy of $\xi_{k,j}$. Conditioned on $\overline{\mathcal{G}}_{k}\doteq\mathcal{G}_{k-1}\cup\sigma\{\boldsymbol{\xi}_{k,M_{1}}\}$, $T_{k}^{x}(\boldsymbol{\xi}_{k,M_{1}},\boldsymbol{\xi}_{k,M_{2}})$ is the sum of i.i.d. centered random variables. In what follows we will often omit the arguments $\boldsymbol{\xi}_{k,M_{1}}$ and/or $\boldsymbol{\xi}_{k,M_{2}}$ from the notations of functions $T_{k}^{x}$. Using representation (24),
@@
To estimate the r.h.s. of the previous inequality we again apply the empirical process method. We first note that for any $x,x^{\prime}\in\mathsf{X}$,
@@
For notational simplicity we will omit dependencies on $k$ and $\boldsymbol{\xi}_{M_{1}}$, and simply write $T^{x}(\cdot)=T_{k}^{x}(\boldsymbol{\xi}_{M_{1}},\cdot)$. Note that the functions in $\mathcal{T}_{k,\boldsymbol{\xi}_{M_{1}}}$ are Lipshitz w.r.t. the uniform metric
@@
We denote the r.h.s. of this inequality by $R_{k}^{\star}$. Clearly, $R_{k}^{\star}$ is $\overline{\mathcal{G}}_{k}$ – measurable function (recall that $\overline{\mathcal{G}}_{k}=\mathcal{G}_{k-1}\cup\sigma\{\boldsymbol{\xi}_{k,M_{1}}\}$). We may conclude that $\operatorname{diam}(\mathcal{T}_{k,\boldsymbol{\xi}_{M_{1}}})\leq 2R_{k}^{\star}$. Furthermore, by (25) its covering number can be bounded as
@@
Applying the tower property we get
$\mathsf{E}[\sup_{x\in\mathsf{X}}|T_{k}^{x}(\boldsymbol{\xi}_{k,M_{2}})|^{2}]\leq\mathsf{E}[\mathsf{E}[\sup_{x\in\mathsf{X}}|T_{k}^{x}(\boldsymbol{\xi}_{k,M_{2}})|^{2}|\overline{\mathcal{G}}_{k}]]$
Using the Dudley integral inequality, e.g. (Vershynin, 2018, Theorem 8.1.6) and assumption A5,
@@
$\lesssim$ $\mathsf{E}[|T^{x_{0}}(\boldsymbol{\xi}_{k,M_{2}})|^{2}|\overline{\mathcal{G}}_{k}]$
$+$ $\frac{1}{M_{2}}\Bigl{\{}L_{T}\sqrt{C_{\mathsf{X},\mathsf{A}}}\int_{0}^{2R_{k}^{\star}/L_{T}}\sqrt{\log(1+1/\varepsilon)}\mathrm{d}\varepsilon+R_{k}^{\star}\Bigr{\}}^{2},$
where $x_{0}\in\mathsf{X}$ is some fixed point. To estimate the first term in the r.h.s. of the previous inequality we apply Hoeffding’s inequality. We obtain

### v5

@@ 8 Conclusion and Future Work
In this work, we propose a new approach towards model-free evaluation of the agent’s policies in RL, based on upper solutions to the Bellman optimality equation (1). To the best of our knowledge, UVIP is the first procedure that allows us to construct non-asymptotic confidence intervals for the optimal value function $V^{\star}$ based on the value function corresponding to an arbitrary policy $\pi$. In our analysis, we consider only infinite-horizon MDPs and assume that sampling from the conditional distribution ${\rm P}^{a}(\cdot|x)$ is feasible for any $x\in\mathsf{X}$ and $a\in\mathsf{A}.$ A promising future research direction is to generalize UVIP to the case of finite-horizon MDPs by combining it with the idea of real-time dynamic programming (see Efroni et al. (2019)). Moreover, plain Monte Carlo estimates are not necessarily the most efficient way to estimate the outer expectation in Algorithm $1$. Other stochastic approximation techniques could also be applied to approximate the solution of (2).
Acknowledgments
This work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4E0002 and the agreement with HSE University № 139-15-2025-009. This research was supported in part through computational resources of HPC facilities at HSE University Kostenetskiy et al. (2021).
@@ Appendix A Proof of the Main Results
Throughout this section we will use additional notation. Let $\psi_{2}(x)=e^{x^{2}}-1$, $x\in\mathbb{R}$. For r.v. $\eta$ we denote $\|\eta\|_{\psi_{2}}\doteq\inf\{t>0:\mathsf{E}\left[\exp\{\eta^{2}/t^{2}\}\right]\leq 2\}$ the Orlicz 2-norm. We say that $\eta$ is a sub-Gaussian random variable if $\|\eta\|_{\psi_{2}}<\infty$. In particular, this implies that for some constants $C,c>0$, $\mathrm{P}(|\eta|\geq t)\leq 2\exp\{-ct^{2}/\|\eta\|_{\psi_{2}}^{2}\}$ and $\mathsf{E}^{1/p}[|\eta|^{p}]\leq C\sqrt{p}\|\eta\|_{\psi_{2}}$ for all $p\geq 1$. Consider a random process $(X_{t})_{t\in T}$ on a metric space $(T,\mathsf{d})$. We say that the process has sub-Gaussian increments if there exists $K\geq 0$ such that
@@
Proof
We apply empirical process methods. To simplify notation, we denote
@@
that is, $Z(x,a)$ is a random process on the metric space $(\mathsf{X}\times\mathsf{A},\rho)$. Below we show that the process $Z(x,a)$ has sub-Gaussian increments. To show this, let us introduce for $\ell\in[M]$
$Z_{\ell}\doteq[V^{\pi}(\psi(x,a,\xi_{\ell}))-\mathsf{E}V^{\pi}(\psi(x,a,\xi_{\ell}))]-[V^{\pi}(\psi(x^{\prime},a^{\prime},\xi_{\ell}))-\mathsf{E}V^{\pi}(\psi(x^{\prime},a^{\prime},\xi_{\ell}))]\,.$
@@
that is, $Z_{\ell}$ is a sub-Gaussian r.v. for any $\ell\in[M]$. Since $Z(x,a)-Z(x^{\prime},a^{\prime})=M^{-1/2}\sum_{\ell=1}^{M}Z_{\ell}$ is a sum of independent sub-Gaussian r.v., we may apply (Vershynin, 2018, Proposition 2.6.1 and Eq. (2.16)) to obtain that $Z(x,a)$ has sub-Gaussian increments with parameter $K\asymp L_{\pi}$. Fix some $(x_{0},a_{0})\in\mathsf{X}\times\mathsf{A}$. By the triangle inequality,
@@
By Dudley’s integral inequality, e.g. (Vershynin, 2018, Theorem 8.1.6), for any $\delta\in(0,1)$,
$\sup_{(x,a),(x^{\prime},a^{\prime})\in\mathsf{X}\times\mathsf{A}}|Z(x,a)-Z(x^{\prime},a^{\prime})|\lesssim L_{\pi}\bigl[I_{\mathcal{D}}+\mathsf{D}\sqrt{\log(2/\delta)}\bigr]\,$
@@
holds with probability $1-\delta$. The last two inequalities and (15) imply the statement.
@@
$\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)=\frac{1}{M_{2}}\sum_{j=M_{1}+1}^{M_{1}+M_{2}}\max_{a}\left\{r^{a}(x)+\gamma\Bigl(\widehat{V}_{k}^{{\rm{up}}}(Y_{j}^{x,a})-V^{\pi}(Y_{j}^{x,a})+\frac{1}{M_{1}}\sum_{\ell=1}^{M_{1}}V^{\pi}(Y_{\ell}^{x,a})\Bigr)\right\}.$
@@
Using that $\gamma L_{\psi}<1$, the maximal Lipschitz constant of $\widetilde{V}_{k}^{{\rm{up},\pi}}(x),k\in\mathbb{N}$ is uniformly bounded by
@@
$\begin{split}&\mathsf{E}^{1/p}[|R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})-V^{*}(x)|^{p}]\leq\\ &\mathsf{E}^{1/p}\left[\left|\max_{a}\left\{r^{a}(x)+\gamma\Bigl(\widehat{V}_{k}^{{\rm{up}}}(Y_{j}^{x,a})-V^{\pi}(Y_{j}^{x,a})+M_{1}^{-1}\sum\nolimits_{\ell=1}^{M_{1}}V^{\pi}(Y_{\ell}^{x,a})\Bigr)\right\}\right.\right.-\\ &\qquad\left.\left.\max_{a}\left\{r^{a}(x)+\gamma(V^{\star}(Y^{x,a})-V^{\star}(Y^{x,a})+{\rm P}^{a}V^{\star}(x)\right\}\right|^{p}\right].\end{split}$
Hence, with Minkowski’s inequality and $|{\rm P}^{a}V^{\star}(x)-\mathsf{E}V^{\pi}(\psi(x,a,\cdot))|\leq\|V^{\pi}-V^{*}\|_{\mathsf{X}}$, we get
$\begin{split}&\mathsf{E}^{1/p}[|R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})-V^{*}(x)|^{p}]\leq\gamma{\mathcal{M}}_{k}+2\gamma\left\|V^{\pi}-V^{{\rm*}}\right\|_{\mathsf{X}}\\ &\qquad\qquad\qquad+\gamma\mathsf{E}^{1/p}\Bigl[\Bigl\|M_{1}^{-1}\sum\nolimits_{\ell=1}^{M_{1}}[V^{\pi}(\psi(\cdot,\cdot,\xi_{k,\ell}))-\mathsf{E}V^{\pi}(\psi(\cdot,\cdot,\xi_{k,\ell}))]\Bigr\|_{\mathsf{X}\times\mathcal{A}}^{p}\Bigr].\end{split}$ (19)
To analyze the last term we use empirical process methods. By Proposition A.1, we get
@@
Furthermore, with (7) we construct a Lipschitz interpolant $\widehat{V}_{k+1}^{{\rm{up}}}$ such that
@@
Proof(Proof of Corollary 6.1)
Consider $\rho((x,a),(x^{\prime},a^{\prime}))=\mathbb{I}_{\{(x,a)\neq(x^{\prime},a^{\prime})\}}$ and $\mathsf{X}_{N}=\mathsf{X}$, that is, we bypass the approximation step. Then $\mathsf{D}=1$, $I_{\mathcal{D}}\lesssim\sqrt{\log(|\mathsf{X}||\mathsf{A}|)}$, $\rho(\mathsf{X}_{N},\mathsf{X})=0$, and $r^{a}(\cdot)$ is Lipschitz w.r.t. $\rho_{\mathsf{X}}$ with $L_{\max}\leq R_{\max}$. Moreover, one can take $L_{\psi}=1$ and $L_{\pi}=R_{\max}/(1-\gamma)$ in Assumption A4. Hence, A1 – A4 are valid and one may apply Theorem 6.1. Bound (21) in this case writes as
@@
We use the definition of $\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)$ and $R_{k}^{x}(\xi_{k,j};\xi_{k,1},\ldots,\xi_{k,M_{1}})$ from Theorem 6.1. To simplify notation, we denote $\boldsymbol{\xi}_{k,M_{1}}=(\xi_{k,1},\ldots,\xi_{k,M_{1}})$ and $\boldsymbol{\xi}_{k,M_{2}}=(\xi_{k,M_{1}+1},\ldots,\xi_{k,M_{1}+M_{2}})$. In this notation $\boldsymbol{\xi}_{k}=(\boldsymbol{\xi}_{k,M_{1}},\boldsymbol{\xi}_{k,M_{2}})$. Recall that, by construction, $\widetilde{V}_{k+1}^{{\rm{up},\pi}}(x)$ can be evaluated only at the points $x\in\{\mathsf{x}_{1},\dots,\mathsf{x}_{N}\}$. By definition,
@@
where $\xi$ is an i.i.d. copy of $\xi_{k,j}$. Conditioned on $\overline{\mathcal{G}}_{k}\doteq\mathcal{G}_{k-1}\cup\sigma\{\boldsymbol{\xi}_{k,M_{1}}\}$, $T_{k}^{x}(\boldsymbol{\xi}_{k,M_{1}},\boldsymbol{\xi}_{k,M_{2}})$ is the sum of i.i.d. centered random variables. In what follows, we will often omit the arguments $\boldsymbol{\xi}_{k,M_{1}}$ and/or $\boldsymbol{\xi}_{k,M_{2}}$ from the notation of functions $T_{k}^{x}$. Using representation (24),
@@
To estimate the right-hand side of the previous inequality we again apply the empirical process method. We first note that for any $x,x^{\prime}\in\mathsf{X}$,
@@
For notational simplicity we will omit dependencies on $k$ and $\boldsymbol{\xi}_{M_{1}}$, and simply write $T^{x}(\cdot)=T_{k}^{x}(\boldsymbol{\xi}_{M_{1}},\cdot)$. Note that the functions in $\mathcal{T}_{k,\boldsymbol{\xi}_{M_{1}}}$ are Lipschitz w.r.t. the uniform metric
@@
We denote the right-hand side of this inequality by $R_{k}^{\star}$. Clearly, $R_{k}^{\star}$ is an $\overline{\mathcal{G}}_{k}$-measurable function (recall that $\overline{\mathcal{G}}_{k}=\mathcal{G}_{k-1}\cup\sigma\{\boldsymbol{\xi}_{k,M_{1}}\}$). We may conclude that $\operatorname{diam}(\mathcal{T}_{k,\boldsymbol{\xi}_{M_{1}}})\leq 2R_{k}^{\star}$. Furthermore, by (25), its covering number can be bounded as
@@
Applying the tower property, we get
$\mathsf{E}[\sup_{x\in\mathsf{X}}|T_{k}^{x}(\boldsymbol{\xi}_{k,M_{2}})|^{2}]\leq\mathsf{E}[\mathsf{E}[\sup_{x\in\mathsf{X}}|T_{k}^{x}(\boldsymbol{\xi}_{k,M_{2}})|^{2}|\overline{\mathcal{G}}_{k}]].$
Using Dudley’s integral inequality, e.g. (Vershynin, 2018, Theorem 8.1.6) and assumption A5,
@@
$\qquad\qquad\qquad\lesssim\mathsf{E}[|T^{x_{0}}(\boldsymbol{\xi}_{k,M_{2}})|^{2}|\overline{\mathcal{G}}_{k}]+\frac{1}{M_{2}}\Bigl\{L_{T}\sqrt{C_{\mathsf{X},\mathsf{A}}}\int_{0}^{2R_{k}^{\star}/L_{T}}\sqrt{\log(1+1/\varepsilon)}\mathrm{d}\varepsilon+R_{k}^{\star}\Bigr\}^{2},$
where $x_{0}\in\mathsf{X}$ is some fixed point. To estimate the first term in the right-hand side of the previous inequality we apply Hoeffding’s inequality. We obtain